OpenStudy (jkbo):

please help. find the arithmetic mean given a^n-l=2 a^n+1=8. thank you

OpenStudy (anonymous):

\[ a_{n-1}=2,\;a_{n+1}=8 \]Is this it?

OpenStudy (jkbo):

yes

OpenStudy (anonymous):

The arithmetic mean is just the sum of the terms divided by the number of terms.

OpenStudy (anonymous):

Do they really want the man, or do they want the sequence? I'm find the question quite strange.

OpenStudy (anonymous):

mean^

OpenStudy (jkbo):

the mean

OpenStudy (jkbo):

i m not sure how to do it though

OpenStudy (anonymous):

Me neither, not because I don't know how, but I don't understand what they want the mean of.

OpenStudy (anonymous):

The mean of 8 and 2 is (8+2)/2 = 5

OpenStudy (anonymous):

But I get the feeling that isn't what they want.

OpenStudy (jkbo):

i think it is its just the arithmetic mean.